You are given the following information about a study in a journal article that
used a One way ANOVA, F(8, 211). You would say there were ____ total
participants in this study.
Answer #1

df of the numerator = 8

df of the denominator = 211

Total participants = 211 + 8 = 219 + 1 = 220 participants

Please give me a thumbs-up if this helps you out. Thank you!